Name of the course

NC-vvv-yyyy 10. Oktober 2000

Title

Report

Author1  Author2  etc.
author1@cosy.sbg.ac.at  author2@cosy.sbg.ac.at  etc.

Academic Supervisor Mag. Dipl.-Ing. Schwaiger Roland

Department of Computer Science University of Salzburg

Correspondence to:

Universität Salzburg Institut für Computerwissenschaften und Systemanalyse Jakob–Haringer–Straße 2 A–5020 Salzburg Austria

Title

Author1 and Author2 and etc.

Department of Computer Science University of Salzburg A–5020 Salzburg, Austria

{author1}, {author2}, {etc.}@cosy.sbg.ac.at

Zusammenfassung

– Length: 100 - 150 words. Purpose: Give the reader an overview of your work, i.e. your approach and the results you achieved –